### v1

# The JARVIS Infrastructure is All You Need for Materials Design

Abstract
Joint Automated Repository for Various Integrated Simulations (JARVIS) is a comprehensive infrastructure offering databases, tools, tutorials, and benchmarks for multiscale, multimodal, forward, and inverse materials design. Emphasizing open access principles and reproducibility, it integrates theoretical and experimental methodologies such as density functional theory, quantum Monte Carlo, tight-binding, classical force fields, and machine-learning approaches-including fingerprinting, graph neural networks, and transformer models. Its experimental data collection spans cryogenics, microscopy, and diffraction, covering materials like metals, semiconductors, insulators, superconductors, carbon capture systems, high-strength compounds, and low-dimensional materials, heterostructures and defects. JARVIS disseminates resources via open datasets, web applications, executable scripts, and peer-reviewed publications, ensuring broad accessibility and reproducibility. Widely adopted worldwide, it has facilitated millions of data and tool downloads. By unifying diverse methods and data under one platform, JARVIS drives both fundamental discoveries and real-world innovations, advancing conventional and data-driven materials design.
[…]
Material Measurement Laboratory, National Institute of Standards and Technology, Gaithersburg, MD 20899, USA
1 Introduction
Materials design has long been a cornerstone of technological progress, underpinning advancements in energy, computing, healthcare, and manufacturing [1]. Traditionally, materials discovery followed an empirical, trial-and-error approach, often requiring decades to identify and optimize materials for specific applications. The advent of computational materials science, particularly with the rise of density functional theory (DFT), molecular dynamics (MD)/force-field (FF) and machine learning (ML) has significantly accelerated the pace of discovery [2, 3, 4]. However, the growing complexity of materials problems, the vast chemical and structural search space, and the increasing demand for high-performance materials necessitate more sophisticated, integrated frameworks that can streamline the entire process from fundamental property prediction to real-world implementation.
An integrated materials design infrastructure bridges the gap between theoretical models, computational simulations, and experimental validation, ensuring that materials research is reproducible, scalable, and accessible. Moreover, multiscale modeling [5] capabilities available in these infrastructures enable realistic materials design. Such frameworks incorporate multimodal datasets, automated workflows, and ML models, enabling both forward design (predicting properties from structures) [6, 7, 8, 9] and inverse design (identifying structures with desired properties) [10, 11, 12, 13]. Forward design remains challenging due to the need for high-throughput calculations across multiple material classes, while inverse design requires advanced optimization techniques, generative models, and robust experimental validation.
One of the critical challenges in forward materials design is the computational cost associated with methods such as first-principles simulations. While DFT and beyond methods are widely used for predicting electronic and structural properties, it is limited by system size and accuracy trade-offs, particularly for strongly correlated materials and high-temperature phases [14]. In classical approaches, empirical force fields enable large-scale simulations but often lack transferability across different materials and conditions [15]. More recently, machine learning property prediction models and machine learning force fields (MLFFs) have emerged as a promising alternative, offering near-DFT accuracy at significantly lower computational cost, yet challenges remain in ensuring generalizability and robustness [16, 17]. For experimental methods, availability of standardized benchmarks, integration of analytical and computational methods at various lengths and time scales and framework to enhance reproducibility is still a critical challenge [18].
Inverse design, on the other hand, relies heavily on data-driven methodologies, requiring extensive databases of materials properties and efficient search algorithms [13, 19]. Traditional approaches, such as genetic algorithms and evolutionary strategies, have been complemented by deep learning models, which can generate candidate materials based on target functionalities [20]. However, a major bottleneck is the integration of experimental validation with computational predictions, as synthetic feasibility and stability assessments are often overlooked in algorithmic designs [21]. Additionally, the lack of standardized benchmarking protocols makes it difficult to compare different methods and validate new materials against experimental findings [22].
Given these challenges, an effective materials design infrastructure must not only provide computational tools but also enable reproducibility, findability, accessibility, interoperability, and reusability (FAIR) [23] data practices, and seamless integration of theoretical, computational, and experimental methods. The development of platforms such as Joint Automated Repository for Various Integrated Simulations (JARVIS), which integrates quantum and classical calculations, machine learning models, and experimental datasets, addresses these gaps by offering a unified ecosystem for forward and inverse materials design. By standardizing workflows, automating simulations, and facilitating community-driven data sharing, such infrastructures pave the way for accelerated discoveries and technological breakthroughs in materials science.
The JARVIS infrastructure [4, 24] distinguishes itself from other materials design frameworks by offering a multimodal, scalable, and reproducible ecosystem that integrates first-principles calculations, machine learning models, experimental datasets, and web-based tools into a unified platform. Databases such as Automatic FLOW for Materials Discovery (AFLOW) [25], Open Quantum Materials Database (OQMD) [10], Novel Materials Discovery (NOMAD) [26], Materials Cloud [27] and open catalyst project (OC20) [28] focus primarily on density functional theory (DFT)-based materials screening. These initiatives have fostered a materials-by-design paradigm, where advanced computation, automation, and artificial intelligence (AI) accelerate the transition from fundamental discoveries to industrial applications for next-generation materials innovation. JARVIS extends its capabilities beyond DFT calculations to include classical force-field, machine learning force fields (MLFF), beyond DFT methods such as quantum Monte Carlo and dynamical mean field theory (DMFT). It also covers all major classes of materials such as superconductors, solar cells, topological materials, low dimensional, catalytic, high-strength, thermoelectric, piezoelectric, dielectric discovery, quantum computing algorithms, and experimental validation datasets, making it a truly multimodal and a multiscale infrastructure. Another key differentiator is reproducibility, which is central to JARVIS through the use of open-access datasets, FAIR-compliant data sharing, and the distribution of workflows via interactive web apps and notebooks, benchmarking platform such as JARVIS-Leaderboard ensuring that users can verify and extend existing research [29]. Additionally, JARVIS is designed for scalability, providing not only extensive high-throughput datasets but also machine learning models capable of accelerating materials discovery at significantly reduced computational cost compared to traditional methods [30].
Unlike various other platforms, which often focus on a single domain (e.g., semiconductors or metals), JARVIS covers a broad range of materials, including superconductors, carbon capture materials, and low-dimensional heterostructures, facilitating both forward and inverse design approaches. Its seamless integration of theoretical and experimental techniques-spanning quantum Monte Carlo, classical force fields, microscopy, and diffraction databases ensures that the infrastructure is not only predictive but also experimentally validated, bridging the gap between computational discoveries and real-world applications. With thousands of users, millions of dataset downloads, and growing adoption across academia, industry, and government research, JARVIS has established itself as a comprehensive infrastructure that accelerates materials innovation through automation, interoperability, and AI-driven insights. In this article, we briefly discuss some of the components, impacts, and outreach of the JARVIS infrastructure. For more detailed information, we refer to our latest references such as [4, 24].
2 Overview of the JARVIS Infrastructure
Figure 1: Core Components of the NIST JARVIS Infrastructure. This schematic highlights key elements of JARVIS, including databases for structural, electronic, mechanical, and topological properties, tools for AI/ML modeling, simulations, and benchmarking, tutorials for educational outreach, benchmarking standards for validation against experiments, and community engagement through workshops and collaborations, fostering reproducible materials research.
2.1 Core Components
JARVIS is a comprehensive infrastructure to accelerate materials discovery and design through the integration of advanced computational tools and extensive datasets. At its core, JARVIS encompasses several key components: 1) databases, 2) software tools, 3) interactive webapps, 4) a large number of benchmarks, 5) step by step tutorials and 6) various events to promote materials design knowledge among students and researchers as shown in Fig. 1. A brief description of each component is given below.
2.1.1 Databases
JARVIS hosts a suite of materials databases that provide detailed information on various material properties. These databases consist of its own resources such as JARVIS-DFT as well as external datasets such as Alexandria DB [31] in a uniform format to allow seamless integration as well as comparative analysis. The aggregate of such databases can be up to 6 million materials and 10 million properties. These databases have been downloaded close to 2 million times indicating a wide adaptability by the materials research community. Note that there could be overlap among various databases. To name a few of these, the JARVIS-DFT database contains density functional theory calculations for over 80,000 materials, offering data on structural, electronic, optical, mechanical, phonon, and topological properties. Additionally, the JARVIS-FF database includes classical force-field calculations for approximately 2,000 materials, facilitating the study of interatomic interactions and molecular dynamics simulations. These databases are continuously updated to incorporate new materials and properties, ensuring a comprehensive and current resource for researchers. A list of currently available databases are available at: https://pages.nist.gov/jarvis/databases/. While these tutorials are static snapshots of datasets, a much more user friendly and interactive list of databases is available at: https://jarvis.nist.gov/#Databases.
2.1.2 Tools
To support materials design, JARVIS provides a range of software tools that allow setting up and analysis of materials design tasks, leveraging artificial intelligence (AI) and machine learning (ML) models, as well as simulation frameworks. For instance, JARVIS-Tools allow integrating with more than 15 software types such as Vienna Ab initio Simulation Package (VASP) [32, 33], Quantum Espresso (QE) [34], Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [15], etc., automatic curation and analysis of results, converting information to webpages, providing links to databases mentioned above etc. Similarly, the Atomistic Line Graph Neural Network (ALIGNN) property predictor [35] utilizes graph neural networks to rapidly predict material properties, while the ALIGNN Force-field tool [36] enables fast structure optimization. All of these tools are open source, available on GitHub and with reasonable documentation to allow both new comers as well as experienced users to utilize resources available in JARVIS. There are hundreds of GitHub stars, and thousands to millions of PyPi downloads of these software tools indicating profound utilization of these novel resources. These tools are designed to streamline the materials design process, thereby reducing the reliance on time-consuming methods such as first-principles calculations. Currently, there are more than 10 such tools available through JARVIS. A collection of such tools can be found at the https://github.com/knc6
2.1.3 Tutorials
Recognizing the importance of education and accessibility, JARVIS offers a collection of Jupyter, Google Colab and Serverless Materials Design (SLMat) [37] notebooks that serve as tutorials for various methods in materials design. These resources cover topics such as electronic structure calculations, force-field development, AI applications, and quantum computing methods. By providing these educational materials, JARVIS aims to equip researchers and students with the necessary skills to effectively utilize its tools and databases in their work. One of the major components of tutorials is the JARVIS-Tools-Notebooks framework https://github.com/JARVIS-Materials-Design/jarvis-tools-notebooks available on GitHub. Currently, there are more than 100 notebooks available in this repository. Another set of documented tutorials are available at https://pages.nist.gov/jarvis/tutorials/ .
2.1.4 Benchmarking Standards
To ensure the reliability and reproducibility of materials design methods, JARVIS has developed the JARVIS-Leaderboard https://pages.nist.gov/jarvis_leaderboard/, an open-source, community-driven platform that facilitates benchmarking across various categories, including AI, electronic structure, force-fields, quantum computation, and experimental methods. This platform allows users to set up custom benchmarks and contribute datasets, code, and metadata, promoting transparency and standardization in materials research [29]. Currently, there are more than 8 million data points and 1300 benchmarks in the leaderboard and it is continuously expanding.
2.1.5 Outreach and Community Engagement
JARVIS is committed to fostering a collaborative scientific community through various outreach initiatives. The platform hosts workshops, webinars, and collaborative projects to engage with researchers, educators, and industry professionals. These efforts aim to disseminate knowledge, gather feedback, and promote the adoption of JARVIS tools and databases in the broader materials science community. Some of these events include JARVIS-School, Artificial Intelligence for Materials Science (AIMS) and Quantum Matters in Material Sciences (QMMS) workshops. Currently, we have hosted more than 4 AIMS, 3 QMMS and 10 JARVIS-Schools. Most of the video recordings are available on our website https://jarvis.nist.gov/events/.
Collectively, these components establish JARVIS as a robust and versatile infrastructure that supports the entire materials design lifecycle, from data generation and analysis to education and community engagement.
2.2 Adhering to FAIR Principles
The JARVIS infrastructure is committed to the FAIR principles-Findability, Accessibility, Interoperability, and Reproducibility-to enhance data management and utilization in materials science.
2.2.1 Findability
Findability is achieved through comprehensive metadata and indexing, enabling users to efficiently locate datasets and tools within the JARVIS platform. The infrastructure’s databases, such as JARVIS-DFT, are meticulously organized, allowing researchers to search for materials based on various properties and criteria.
2.2.2 Accessibility
Accessibility is ensured by providing open access to JARVIS datasets and tools. Researchers can freely retrieve and utilize data without restrictive barriers, promoting widespread use and collaboration. The platform’s user-friendly web interface and Application Programming Interface (API)s facilitate seamless data access and integration into external workflows.
2.2.3 Interoperability
Interoperability is supported through the use of standardized data formats and protocols for datasets, analysis tools, web design, benchmarking procedures, enabling integration with other databases and computational tools. JARVIS employs common file formats and adheres to community standards, ensuring compatibility and facilitating data exchange across various platforms.
2.2.4 Reusability
Reusability is a cornerstone of JARVIS, achieved by providing detailed documentation, workflows, and tutorials. The platform offers Jupyter and Google Colab notebooks that guide users through data analysis and simulation processes, ensuring that studies can be replicated and validated by others.
By adhering to these FAIR principles, JARVIS enhances the efficiency, transparency, and collaborative potential of materials research, fostering an environment where data and tools are readily available and usable by the global scientific community.
3 Theoretical and Experimental Method Integration
Figure 2: Evolution of Different Projects in the NIST JARVIS Infrastructure (2017–2025). This figure illustrates the timeline and development of various JARVIS initiatives, including JARVIS-DFT, JARVIS-FF, JARVIS-ML, JARVIS-Exp, and other related projects within the NIST JARVIS ecosystem. The steady expansion from 2017 to 2025 highlights the integration of density functional theory (DFT) databases, machine learning models, force field developments, and experimental data repositories. The evolution reflects increasing computational capabilities, data-driven approaches, and interdisciplinary collaborations aimed at advancing materials design and discovery.
Several theoretical and experimental methods have been developed, utilized, and integrated within JARVIS over the course of 7+ years as shown in Fig. 2. It shows the timeline and development of various JARVIS initiatives, including JARVIS-DFT, JARVIS-FF, JARVIS-ML, JARVIS-Exp, ALIGNN, Atomistic Generative Pretrained Transformer (AtomGPT) [38] and other related projects within the JARVIS ecosystem. The steady expansion from 2017 to 2025 highlights the integration of density functional theory (DFT) databases, machine learning models, force field developments, and experimental data repositories. The evolution reflects increasing computational capabilities, data-driven approaches, and interdisciplinary collaborations aimed at advancing materials design and discovery. We can categorize these methods into quantum mechanical, machine learning, classical and experimental categories. A brief description of these components is given below.
3.1 Quantum Mechanical Methods
The JARVIS infrastructure integrates a range of quantum mechanical methods to provide comprehensive insights into material properties.
3.1.1 Density Functional Theory and Tight-binding
Because of its ab-inito nature and wide usability, DFT is a cornerstone of JARVIS. Density Functional Theory (DFT) is a computational quantum mechanical modeling method employed to investigate the electronic structure of many-body systems. The JARVIS-DFT database encompasses a vast array of materials, offering detailed information on structural, electronic, optical, and mechanical properties [39, 40, 41, 4], thereby serving as a valuable resource for researchers in materials science. The DFT databases in JARVIS are JARVIS-DFT which is mainly based on VASP calculations and JARVIS-QETB databases which is based on Quantum Espresso (QE) based calculations. DFT has various levels of fidelity such as local-density approximations (LDA), generalized gradient approximation (GGA), metaGGA, van der Waals (vdW) corrected functionals such as OptB88vdW and many body based methods which have various levels of computational cost. JARVIS encompasses most of these techniques in JARVIS-DFT, JARVIS-QETB as well in the JARVIS-Leaderboard framework. JARVIS-QETB was used to develop a universal set of tight-binding parameters across the periodic table [42]. It includes both two-body and three-body effective interaction terms, plus self-consistent charge transfer, enabling the model to work for metallic, covalent, and ionic bonds with the same parameter set.
3.1.2 Dynamical Mean Field Theory
In addition to DFT, JARVIS incorporates more advanced quantum mechanical approaches to address the limitations of standard DFT, particularly in dealing with strongly correlated electron systems [43, 44]. One such method is Dynamical Mean Field Theory (DMFT), which provides a non-perturbative treatment of local interactions and is effective in capturing the physics of correlated materials. By integrating DMFT, JARVIS enhances its capability to predict electronic properties in materials where electron correlation plays a significant role.
3.1.3 Quantum Monte Carlo
Furthermore, JARVIS employs Quantum Monte Carlo (QMC) [45, 46] methods, which are stochastic approaches used to solve the Schrödinger equation with high accuracy. QMC is particularly useful for benchmarking DFT results and providing insights into systems where DFT may be inadequate. The inclusion of QMC calculations within JARVIS allows for more reliable predictions of material properties, especially in cases involving complex electronic interactions.
3.1.4 Quantum Computing
Recognizing the emerging potential of quantum computing in materials science, JARVIS also explores quantum computing algorithms for simulating material properties [47, 44]. By interfacing with quantum computing frameworks such as Qiskit, JARVIS aims to leverage quantum algorithms to solve problems that are computationally intensive for classical computers, such as the simulation of strongly correlated systems and the calculation of electronic spectra. This integration positions JARVIS at the forefront of incorporating cutting-edge computational techniques in materials research.
By encompassing these diverse quantum mechanical methods, JARVIS provides a robust and versatile platform for the accurate prediction and analysis of material properties, catering to a wide range of research needs in the field of materials science.
3.2 Machine Learning and Data-Driven Approaches
3.2.1 Fingerprinting Techniques
The JARVIS infrastructure leverages advanced machine learning and data-driven approaches to enhance materials discovery and design. A key aspect of this strategy involves the use of fingerprinting techniques, which generate unique representations of materials’ structural and compositional features such as element fraction descriptors, average chemical descriptors and classical force-field inspired descriptors (CFID) descriptors [48]. JARVIS is also integrated within other fingerprinting tools such as MatMiner [49] to allow various other schemes of fingerprints which can then be compared for instance in the JARVIS-Leaderboard [29] for accuracy and speed. These fingerprints serve as inputs for machine learning models. These models have recently been analyzed for their uncertainty quantification applications as well [50]. Very often, uncertainty quantification metrics can be more important than typical machine learning model metrics while implementing realistic materials design.
3.2.2 Graph Neural Networks
JARVIS employs Graph Neural Networks (GNNs) to model the complex relationships within materials. Such GNNs have been used for atomic structure and atomistic image analysis in ALIGNN [35], ALIGNN-FF [36] and AtomVision [51] frameworks. They are becoming an inevitable part of almost all materials design processes. GNNs are adept at handling graph-structured data, making them well-suited for representing materials where atoms are nodes and bonds are edges. The Atomistic Line Graph Neural Network (ALIGNN) is a notable example, which enhances traditional GNNs by incorporating both bond lengths and bond angles into its architecture. This inclusion of angular information has led to improved performance in predicting a wide range of material properties, as demonstrated in studies utilizing the JARVIS-DFT and other external datasets.
3.2.3 Transformer Models
In addition to GNNs, JARVIS explores the application of transformer-based models in materials science. Originally developed for natural language processing( NLP such as in ChemNLP [52]), Transformers have been adapted to handle the complex patterns found in material structures for both forward and inverse materials design such as in AtomGPT [38] and DiffractGPT [53] frameworks. Recent research has introduced models that generate atomic embeddings using Transformer architectures, leading to enhanced predictions of material properties. These models capture long-range dependencies and intricate relationships within the material data, offering a powerful tool for materials informatics.
By integrating these machine learning methodologies-fingerprinting techniques, Graph Neural Networks, and Transformer-based models—JARVIS provides a robust framework for the accelerated discovery and design of new materials.
3.3 Classical Methods and Force Fields
The JARVIS infrastructure integrates classical simulation methods and force fields to complement its quantum mechanical approaches, thereby providing a comprehensive toolkit for materials modeling.
3.3.1 Conventional Force Fields
A fundamental component of JARVIS is Classical Molecular Dynamics (MD), which utilizes empirical interatomic potentials to simulate the behavior of atoms and molecules over time. This method is particularly effective for studying the thermodynamic and kinetic properties of materials at finite temperatures and for modeling large systems that are computationally prohibitive for quantum mechanical methods. JARVIS-FF [54, 55] maintains a repository of classical force fields tailored for various materials, facilitating accurate MD simulations within its framework with Large-scale Atomic/Molecular Massively Parallel Simulator, General Utility Lattice Program (GULP), Assisted Model Building with Energy Refinement (AMBER) and similar MD software packages.
3.3.2 Machine Learning Force Fields
To enhance the accuracy and transferability of force fields, JARVIS incorporates Machine Learning Force Fields (MLFFs) [56] based on ALIGNN. The ALIGNN-FF is seamlessly integrated with an atomic simulation environment (ASE) Calculator to leverage various dynamic tools. These MLFFs are trained on extensive datasets derived from high-fidelity quantum mechanical calculations, enabling them to capture complex potential energy surfaces with high precision. By leveraging machine learning techniques, JARVIS’s MLFFs can predict interatomic forces and energies more efficiently than traditional empirical potentials, thereby expanding the scope and scale of simulations that can be performed. This approach allows for rapid and accurate predictions of material properties, facilitating accelerated materials discovery and design.
By incorporating classical MD and MLFFs, JARVIS offers a robust and flexible infrastructure for the multiscale simulation of materials, accommodating a wide range of systems and properties with varying degrees of complexity.
3.4 Experimental Data Integration
The JARVIS infrastructure emphasizes the integration of experimental data to validate and enhance its theoretical predictions, thereby bridging the gap between computation and real-world observations. A key aspect of this integration involves the incorporation of experimental data for benchmarking materials properties predicted with quantum/classical methods. While most of the benchmarking experimental data comes from previous experiment literature, JARVIS has its own experimental data including microscopy, diffraction and cryogenics experiments. By comparing simulated data with experimental data, JARVIS ensures the accuracy and reliability of its computational models. Furthermore, JARVIS integrates data for inorganic, organic and metal-organic compounds. Some examples include experimental measurements of superconductors [57], nanoparticles [54], topological magnets [58], solar cells [40] , metal organic frameworks [59] etc. This integration allows for the validation of theoretical predictions against experimentally determined crystal structures, enhancing the robustness of the simulations.
By correlating theoretical predictions with experimental observations, JARVIS not only validates its computational approaches but also refines its models to better reflect real-world behaviors. This synergy between theory and experiment is crucial for advancing materials science, as it ensures that computational discoveries are grounded in empirical evidence, thereby accelerating the development of new materials with desired properties. We note that while automating theoretical materials design methods are well developed the same for experiments is still lacking. In the near future, JARVIS will focus on automating and benchmarking every distinct aspect of materials experimental methods, enabling truly autonomous self-driving lab experiments and advancing AI-driven materials design.
4 Coverage Across Material Classes
The JARVIS infrastructure offers extensive coverage across a diverse array of material classes, facilitating advancements in both theoretical understanding and practical applications as shown in Fig. 3. It shows various types of materials classes with an example of each available in JARVIS such as metals, semiconductors, insulators, alloys [40], topological insulators [60, 61], superconductors [62, 57, 63, 64], solar absorbers [41], mechanically hard materials [39], intercalated materials for battery cathodes, mechanically exfoliable 2D, 1D, 0D materials[39, 65], crystalline polymers, metal organic frameworks [59], low infrared active mode, high infrared active mode materials, materials with dielectric, piezoelectric constants, ferroelectrics [66], thermoelectrics [67], 2D ferromagnets [45], anomalous quantum confinement effect materials, Dirac and Weyl semimetals, quantum anomalous hall and spin hall insulators [61], Chern insulators [58], molecules in Computational Chemistry Comparison and Benchmark DataBase (CCCBDB) database [68], proteins in external protein databank (PDB) datasets [69], amorphous materials generated with ALIGNN-FF/DFT [70], gas/solid interface for catalysis [71], solid-solid interfaces for microelectronics [72] and other defect based materials [73]. Unless specified with the database name, the identifier should indicate JARVIS-DFT identifiers. A brief description of some of these classes is given below.
Figure 3: Atomic structure, identifier and chemical formula for various materials classes are provided wherever applicable. Examples include metals, semiconductors, insulators, alloys, topological insulators, superconductors, solar absorbers, mechanically hard materials, intercalated materials for battery cathodes, mechanically exfoliable 2D, 1D, 0D materials, crystalline polymers, metal organic frameworks, low infrared active mode, high infrared active mode materials, materials with dielectric, piezoelectric constants, ferroelectrics, thermoelectrics, 2D ferromagnets, anomalous quantum confinement effect materials (AQCE), Dirac and Weyl semimetals, quantum anomalous hall (QAHI) and spin hall insulators (QSHI), Chern insulators, molecules in CCCBDB database, proteins in external PDB datasets, amorphous materials generated with ALIGNN-FF/DFT, gas/solid interface for catalysis, solid-solid interfaces for microelectronics and other defect based materials.
4.1 Metals, Semiconductors, and Insulators
Based on electronic bandgaps, materials can be classified into metal, semiconductor and insulator categories. JARVIS provides comprehensive datasets on these fundamental material categories, encompassing detailed information on their structural, electronic, and mechanical properties. This data is crucial for applications ranging from electronic device fabrication to structural engineering. For instance, the platform’s resources aid in the selection and optimization of materials for semiconductors, which are integral to modern electronics.
4.2 Superconductors and Quantum Materials
All materials are quantum in nature, but when these effects show dominance at the classical level as well, then they can be utilized for numerous technological applications. Recognizing the significance of quantum materials in advancing technology, JARVIS includes data on superconductors and other quantum materials such as topological insulators, anomalous Hall insulators, Dirac and Weyl semimetals etc. This encompasses properties such as critical temperatures and electronic band structures, which are vital for developing applications like quantum computing and highly efficient power transmission systems.
4.3 Carbon Capture and Sustainable Materials
Addressing environmental challenges, JARVIS integrates data on materials designed for carbon capture and sustainability. This includes information on metal-organic frameworks (MOFs), zeolytes, polymeric membranes and other porous materials that can adsorb carbon dioxide, contributing to efforts in mitigating climate change.
4.4 High-Strength and Structural Materials
For applications requiring materials with exceptional strength and durability, such as in the construction and aerospace industries, JARVIS can be used to rank the strength of materials based on the elastic modulus and derived data. This information assists in the design and selection of materials that meet specific mechanical performance criteria.
4.5 Low-Dimensional Materials
The dimensionality of a material can be defined based on the presence of vdW bonding in 1,2,3 crystallographic directions. JARVIS provides extensive data on 0D, 1D, 2D materials and their properties, which have unique electronic and mechanical properties. This includes information on materials like graphene and transition metal dichalcogenides, which are being explored for next-generation electronic devices and sensors.
4.6 Defects and Interfaces
Understanding defects and interfaces/heterostructures is crucial for tailoring material properties for specific applications. JARVIS includes data on various defect structures and their properties such as vacancies, enabling researchers to predict how these imperfections can influence material behavior. This is particularly important in semiconductors, where defects can significantly impact electronic properties. Similarly, JARVIS provides databases, tools, webapps, and benchmarks on solid-molecule interface (such as for catalysts) and solid-solid interface (such as for heterostructures used in microelectronics) are crucial for materials design.
By encompassing such a wide range of material classes, JARVIS serves as a valuable resource for researchers and engineers, supporting the development of innovative materials and technologies across multiple industries.
5 JARVIS as a Platform for Reproducible Science
The JARVIS serves as a comprehensive platform dedicated to promoting reproducible science in materials research. There are various strategies to enhance reproducibility and descriptions of these strategies are briefly given below.
5.1 Peer-Reviewed Publications and Open-Access Data
JARVIS emphasizes transparency by providing open-access datasets and disseminating findings through peer-reviewed publications. Ideally, each publication in JARVIS should have: 1) fully accessible preprint on arXiv/ChemarXiv etc., 2) code available in GitHub or similar platforms, 3) static data available in Figshare within JARVIS-Tools or similar platforms, 4) benchmarks used for methods available in JARVIS-Leaderboard, 5) WebApp to enhance the applicability of dataset, tools, models generated. This approach ensures that the scientific community can readily access and validate the data, fostering trust and facilitating further research. The platform’s commitment to open science is evident in its extensive databases, such as JARVIS-DFT, which offers a wealth of information on material properties.
5.2 Interactive Web Applications
To enhance user engagement and accessibility, JARVIS offers a suite of web applications that provide intuitive interfaces for data retrieval and analysis. These tools enable researchers to explore datasets, perform simulations, and visualize results without the need for extensive computational resources, thereby lowering the barrier to entry for materials research.
5.3 Scripts and Notebooks for Reproducibility
Recognizing the importance of reproducibility, JARVIS provides a collection of scripts and Jupyter notebooks that detail the methodologies used in data generation and analysis. These resources allow researchers to replicate studies, validate findings, and build upon existing work, ensuring that scientific discoveries are robust and verifiable.
5.4 Benchmarking and Validation
JARVIS has developed the JARVIS-Leaderboard, an open-source, community-driven platform designed to facilitate benchmarking across various materials design methods. This initiative allows for the comparison of computational predictions with experimental datasets, providing a framework for the performance evaluation of different computational approaches. By systematically validating methods against empirical data, JARVIS ensures the reliability and accuracy of its predictive models.
Through these initiatives, JARVIS not only advances materials science but also sets a standard for reproducibility and transparency in computational research.
6 Impact and Community Adoption
Figure 4: Trends in Citations, Software Downloads, and Geospatial Data Visualization. (a) Annual citation count from 2018 to 2025, shows a significant increase in recent years. (b) PyPi downloads for various materials science and AI-related tools, formatted using engineering notation for clarity. (c) Figshare downloads over time, highlighting adoption trends across different years. Note that some of the PyPI and Figshare download metrics may be influenced by intentional or unintentional automation using bots. Therefore, these numbers should be interpreted with caution. (d) A map visualization provides geospatial insights into usage patterns. Darker blue indicates a higher number, while lighter shades represent lower values.
The JARVIS has established itself as a pivotal resource in the materials science community, fostering global engagement and collaboration across academia and industry.
6.1 Global User Base and Downloads
Since its inception in January 2017, JARVIS has garnered a substantial international user base, with over 150,000 users accessing its resources. The papers in JARVIS have been cited over 4000 times. The platform’s extensive databases, encompassing more than 80,000 materials and over a million calculated properties, have been downloaded over a million times, underscoring its widespread adoption and utility in the research community. Examples of impact in terms of paper citations, software and data downloads, and global user base are shown in Fig. 4.
6.2 Industry and Academic Collaborations
JARVIS actively supports real-world research and development by facilitating collaborations between academic institutions and industry partners. By providing open-access data and computational tools, JARVIS enables researchers to accelerate materials discovery and optimization, bridging the gap between theoretical research and practical applications. These collaborations have led to advancements in various fields, including electronics, energy storage, and sustainable materials.
In addition, to promote education and skill development, JARVIS organizes events such as the JARVIS-School, AIMS and QMMS workshops which offer tutorials and hands-on sessions on topics like electronic structure calculations, machine learning applications, and quantum computations as mentioned in the above section of outreach. These programs are designed to introduce participants to open-access databases and tools for materials design, fostering a deeper understanding and practical proficiency in modern computational methods.
6.3 JARVIS in Education and Curriculum Development
Beyond workshops, JARVIS aims to contribute to curriculum development by providing educational resources that integrate into academic programs. The platform’s collection of Jupyter and Google Colab notebooks serves as valuable teaching aids, allowing students to engage with real-world data and computational tools in a classroom setting. This integration enhances the learning experience and prepares students for careers in materials science and engineering.
Through these initiatives, JARVIS not only advances materials research but also plays a crucial role in education and the dissemination of knowledge, supporting the development of the next generation of scientists and engineers.
7 Future Directions and Open Challenges
The JARVIS is poised to advance materials science through several strategic initiatives, while also addressing existing challenges.
7.1 Expanding Materials Coverage
The number of possible materials and types is huge. JARVIS aims to broaden its database in the future by incorporating a wider array of materials, including complex compounds, alloys, and emerging materials like altermagnets, multi-layer heterostructures, etc. This expansion will enhance the platform’s utility for diverse research applications.
7.2 Enhancing AI/ML Capabilities
The integration of advanced artificial intelligence and machine learning techniques is a priority for JARVIS. By developing more sophisticated models, such as graph neural networks and transformer-based architectures, the platform seeks to improve the accuracy and efficiency of property predictions and materials discovery.
7.3 Improving Quantum Computing Integration
Recognizing the potential of quantum computing, JARVIS plans to deepen its integration of quantum algorithms to tackle complex simulations that are challenging for classical computers. This includes exploring quantum algorithms for electronic structure calculations and materials optimization.
7.4 Addressing Scalability and High-Performance Computing Needs
With growing length-time scale requirements, and database growth, ensuring scalability and efficient data management becomes crucial. JARVIS is investing in high-performance computing resources and optimizing its computational workflows to handle large-scale simulations and data analyses effectively.
7.5 Strengthening Experimental-Theoretical Synergy and Autonomous Experiments
Finally, to bridge the gap between theory and experiment, JARVIS is enhancing its integration of experimental data, such as microscopy and spectroscopy results, with computational predictions. This synergy aims to improve the validation of theoretical models and foster a more comprehensive understanding of material behaviors, which are essential for self-driving labs, autonomous experimentation and robotic agent development.
By pursuing these directions, JARVIS is committed to overcoming current challenges and advancing the field of materials science through innovation and collaboration.
8 Conclusions
The JARVIS has significantly advanced materials science by providing an integrated platform that combines electronic structure calculations, machine learning models, and experimental data. This comprehensive approach has facilitated accelerated materials discovery and design, contributing to numerous peer-reviewed publications and the development of open-access datasets. Looking ahead, JARVIS is poised to play a pivotal role in the next decade of materials discovery. By expanding its databases to include emerging materials and enhancing its machine learning capabilities, JARVIS aims to provide more accurate predictions and insights. The integration of quantum computing algorithms and the development of benchmarking platforms like the JARVIS-Leaderboard further position it as a leader in the field. To fully realize these advancements, collaboration with the broader scientific community is essential. Researchers, educators, and industry professionals are encouraged to engage with JARVIS by utilizing its resources, contributing data, and participating in community-driven initiatives. Such collective efforts will enhance the platform’s capabilities and drive innovation in materials science.
9 Data Availability
The JARVIS data is available at https://jarvis.nist.gov/.
10 Acknowledgments
K.C thanks the National Institute of Standards and Technology for funding, computational, and data-management resources. This work was performed with funding from the CHIPS Metrology Program, part of CHIPS for America, National Institute of Standards and Technology, U.S. Department of Commerce. Certain commercial equipment, instruments, software, or materials are identified in this paper in order to specify the experimental procedure adequately. Such identifications are not intended to imply recommendation or endorsement by NIST, nor it is intended to imply that the materials or equipment identified are necessarily the best available for the purpose. This is an Official contribution of the National Institute of Standards and Technology; not subject to copyright in the United States.